# SPIQA: A Dataset for Multimodal Question Answering on Scientific Papers

### v1

@@ Abstract
Seeking answers to questions within long scientific research articles is a crucial area of study that aids readers in quickly addressing their inquiries. However, existing question-answering (QA) datasets based on scientific papers are limited in scale and focus solely on textual content. To address this limitation, we introduce SPIQA (Scientific Paper Image Question Answering), the first large-scale QA dataset specifically designed to interpret complex figures and tables within the context of scientific research articles across various domains of computer science. Leveraging the breadth of expertise and ability of multimodal large language models (MLLMs) to understand figures, we employ automatic and manual curation to create the dataset. We craft an information-seeking task involving multiple images that cover a wide variety of plots, charts, tables, schematic diagrams, and result visualizations. SPIQA comprises 270K questions divided into training, validation, and three different evaluation splits. Through extensive experiments with 12 prominent foundational models, we evaluate the ability of current multimodal systems to comprehend the nuanced aspects of research articles. Additionally, we propose a Chain-of-Thought (CoT) evaluation strategy with in-context retrieval that allows fine-grained, step-by-step assessment and improves model performance. We further explore the upper bounds of performance enhancement with additional textual information, highlighting its promising potential for future research and the dataset’s impact on revolutionizing how we interact with scientific literature. equal technical contribution, work done as a student researcher at Google Research.Corresponding author (vsubhashini@google.com)
@@ 1 Introduction
Numerous datasets have been curated to evaluate the QA abilities of Large Language Models (LLMs) on various documents, including factual documents [80, 60, 21, 81, 31, 12, 65, 67], book chapters [27, 28, 49], news articles [69, 35] and more. However, understanding scientific papers poses unique challenges as the systems must comprehend underlying theoretical implications with domain-specific terminologies and verify claims with experimental results and visualizations. There are also several datasets that focus on the comprehension of standalone science diagrams [26, 23, 17], mathematical figures [43, 82], charts [5, 46, 47], plots [48], and tables [58, 71, 9, 8, 11, 25, 50]. However, simultaneously reasoning over all figures and associated text in a scientific article necessitates both multimodal and long-context capabilities.
In contemporary research, figures and tables with associated captions are crucial to understand the motivation and contributions of a work. Comprehending such multimodal content presents a significant real-world challenge. In this work, we introduce SPIQA (Scientific Paper Image Question Answering), the first large-scale QA dataset specifically designed to interpret complex figures and tables in the context of scientific research articles across various domains of computer science. We also develop three well-designed tasks using the SPIQA dataset to assess growing long-context capabilities of MLLMs: (1) Direct QA with figures and tables, where the systems require to answer questions after seeing all figures and tables from a scientific paper, (2) Direct QA with full paper, where the systems analyze the whole paper text along with figures and tables to answer questions and (3) CoT QA, a chain of thought (CoT) retrieval based QA where systems need to first identify helpful figures and then answer the question, allowing to evaluate the models for fine-grained reasoning and grounding capability. We collect the PDFs and source TeXs of 26K papers from various domains of computer science published in top-tier academic conferences and generate 270K question-answer-rationale triplets focusing on the papers’ figures and tables.We further filter and augment two existing scientific QA datasets, QASA [34] and QASPER [14], to identify questions requiring reasoning over figures, tables, and textual paragraphs. Overall, SPIQA contains one training, one validation, and three evaluations split with varying difficulty levels, allowing us to fine-tune and assess the capabilities of large multimodal systems to understand complex scientific research papers. An illustration of the three proposed tasks with example prompts is presented in Figure 1.
@@ 2 Related Works
Question-answering on long documents is a challenging real-world task that has attracted increasing attention in recent years, following the success of the long-context reasoning ability of LLMs [39, 10, 68, 61, 52, 2, 36]. Though there exist numerous general-domain document QA datasets [80, 60, 21, 81, 31, 12, 65, 67, 74, 59, 79, 84, 73, 69, 35], understanding scientific papers requires domain-specific expertise and reasoning capability, and hence poses a more significant challenge.
@@
Datasets for QA on Scientific Papers. In the early days, cloze-style academic paper QA datasets [56, 54, 77, 66, 57] were automatically constructed by extracting entities and relations and matching them with structure knowledge resources. The questions in such datasets follow a pre-defined format; hence, they are unsuitable for real-world usage where the reader asks detailed open-ended questions [31]. To overcome these issues, PubMedQA [20], BioAsq [29] and QASPER [14] construct corpora of 1K, 3.2K, and 5K human-written questions, respectively. However, the annotators in these datasets only read the abstracts when writing the questions, and hence, most questions are simple and can be answerable in yes/no format or with short extractive spans. ArgSciChat [63] proposes a dataset of argumentive dialogues between scientists on 20 NLP papers. Closer to our work, QASA [34] generates 1798 free-form advanced questions on AI/ML papers where the annotators read the whole paper. However, QASA questions are answerable just from the text paragraphs, neglecting the structured information present in terms of figures and tables.
@@
Datasets for QA on Scientific Diagrams. Solving mathematical problems in a visual context has emerged as a complex reasoning task for MLLMs. Prior attempts, such as GeoQA [7], UniGeo [6], and Geometry3K [44], have exclusively focused on solving geometry-oriented questions. In a different line of research, datasets like DVQA [22], LEAF-QA [5], LEAFQA++ [64], FigureQA [23], PlotQA [48], and ChartQA [46] have been constructed to solve plot and chart-oriented questions. Additionally, there are datasets for QA purely on tabular data, including WTQ [58], TableQA [71], SQA [19], HiTab [11], AIT-QA [25], FetaQA [50], MultiTabQA [53]. More recently, MathVista [43], and MathVerse [82] have integrated different scientific diagrams to develop benchmarks with a wider variety of tasks. However, all existing datasets focus on asking questions about standalone figures or tables. SPIQA bridges this gap by proposing the first scientific QA benchmark that includes questions that require simultaneous reasoning over figures, tables, and textual paragraphs, allowing a more integrated understanding of scientific documents.
@@ 3.1 Collection Guidelines and Task Formulation
  • CoT QA: To assess the step-by-step reasoning abilities of MLLMs, CoT QA requires systems first to identify relevant figures and tables and then answer the question. For simplicity, CoT QA does not involve full paper text. Detailed prompts for each task are provided in the supplement.
@@
3.2 SPIQA: Data Collection, Question Generation and Filtering
Automatic Question Generation and Filtering. After gathering the main text, figures, and tables from research papers, the next step is to generate high-quality question-answer pairs that cover all aspects of the articles. Manually annotating quality questions requires domain expertise and a deep understanding of the research papers, resulting in existing human-annotated datasets being small-scale or solely based on abstracts [70, 20, 29, 14, 34]. To bridge this gap, we automatically generate QAs by leveraging recent advances in powerful multi-modal large language models. We first conducted a pilot study by selecting 30 papers from various domains and experimenting with multiple models. In our approach, we presented one figure or table to the model, along with passages referencing the figure. We then asked the models to generate a question, answer, and rationale that requires a holistic understanding of the figure or table in the context of the paper. We manually verified the quality of the generated QAs by each model using the following criterion: $(i)$ Answering the question would require a complete understanding of the figure and its importance in the paper. $(ii)$ The generated answer is correct and to the point. $(iii)$ The question is neither too trivial nor too specific to the figure or table. Our pilot study resulted in Gemini 1.5 Pro being the best model for generating the QAs, and we generated 270,194 QAs over 25,859 papers using Gemini. In the supplementary section, we show our final QA generation prompt and examples of QAs generated by different models.
After generating QAs, we divide the dataset into three splits: 200 papers for evaluation, 200 for validation, and the remaining 25,459 for training, ensuring each split consists of papers from all domains. Despite the impressive question quality, we perform additional manual filtering on the evaluation set to exclude any inconsequential and incorrect QAs. In addition to the three criteria of the pilot study, the filtering guideline contains the following standards: $(iv)$ If two or more questions from a paper are similar, keep one. $(v)$ If the question is entirely based on the passage, discard the QA. $(vi)$ If the answer is not clear, e.g., the answer says ‘It is hard to answer the question based on the given information‘ or ‘The answer is not evident from the given passage‘, discard the QA. $(vii)$ If the QA includes phrases like ‘Based on the passage,’ modify it because we show all figures and tables to the model at once during evaluation. The user interface (UI) and detailed guidelines for filtering with examples are presented in the supplementary. Since only around 11% of questions are discarded during filtering the evaluation set, we do not perform additional filtering on the train and validation splits. After filtering, the final evaluation set contains 666 questions, and we refer to this split as test-A.
Split # Papers # Ques. # Figures # Tables
Sche. Plots & Charts Vis. Others.
Train 25,459 262,524 44,008 70,041 27,297 6,450 114,728
Val 200 2,085 360 582 173 55 915
test-A 118 666 154 301 131 95 434
test-B 65 228 147 156 133 17 341
test-C 314 493 415 404 26 66 1332
@@ 3.2 SPIQA: Data Collection, Question Generation and Filtering
Additional Evaluation Sets with Human Written QAs. Since the questions and answers in test-A are automatically generated by Gemini, to develop a robust and complete QA benchmark on scientific articles, we curate two additional test sets with human-written QAs by utilizing two existing datasets, QASA [34] and QASPER [14]. However, the human annotators for these datasets mainly considered the paper text. We aim to identify the questions in these datasets that require reasoning with figures or tables for comprehensive answers. For every question, both QASA and QASPER provide evidence sentences from the paper, which are crucial for answering. We parsed these evidence sentences to look for phrases like ‘According to Figure x,’ ‘As shown in Table y,’ ‘Figure z presents,’ and separate such questions, and treat the corresponding figure and tables as evidence. We collect the PDFs for the papers in QASA and QASPER from arXiv, and extract the figures using PDFFigures 2.0. We conduct additional manual filtering to verify that the identified evidence figures help answer the questions. This process resulted in 228 questions from QASA and 493 questions from QASPER, where figures or tables are essential for answering. We refer to these sets generated from QASA and QASPER as test-B and test-C, respectively. The UI for filtering is shown in the supplementary.
@@ 3.3 Dataset Analysis
Splits & Statistics. The main statistics of the collected papers, generated questions, and data splits are presented in Tables 1 and 2. SPIQA encompasses 25,859 computer science papers published in top-tier conferences between 2018-2023, containing 152,487 figures and 117,707 tables. Figure 2 shows the distribution of figures and tables in every paper. We generate a total of 270,194 QAs focusing on the figures and tables along with the main text of the papers. The questions and answers, on average, contain 12.98 and 14.56 words, respectively. Notably, we observe high variances in their lengths - 20.47 for questions and 243.29 for answers - indicating a diverse range of patterns in SPIQA. The training, validation, and test-A splits include papers from every source conference and ensure that questions from the same paper remain in the same split. The test-B and test-C splits are generated from QASA [34] and QASPER [14] and contain human-written QAs. Examples from the dataset, illustrated in Figure 4, highlight two different questions centered on different types of figures. Additionally, we provide a comprehensive comparison of SPIQA with existing scientific QA datasets in the supplementary materials.
@@ 4 LLMLogScore (L3Score): An Improved Metric for Free-form QA
We alleviate the necessity of a predefined scale range and detailed interpretation of every point in the scale by proposing LLMLogScore (L3Score), which directly uses the log-likelihood probabilities generated by an LLM for evaluating the answers. We use GPT-4o [52] as our LLM, show the model the candidate and the ground-truth answers, and ask if the semantic meaning of the candidate and the ground-truth are similar in the context of the question. The LLM is expected to answer yes or no in a single word. We show the exact prompt used for calculating L3Score in the supplementary. Next, instead of considering the final response by the LLM, we look into the top-5333We choose $n=5$ for top-$n$ log probabilities as this is the highest value of $n$ returned by the OpenAI API. log probabilities and corresponding tokens, and we define the L3Score metric as follows:
@@
Figure 4 shows QA evaluations where L3Score is more appropriate than ROUGE and BERTScore. We provide additional examples comparing L3Score with other LLM-based metrics in supplementary.

### v3

@@ Abstract
Seeking answers to questions within long scientific research articles is a crucial area of study that aids readers in quickly addressing their inquiries. However, existing question-answering (QA) datasets based on scientific papers are limited in scale and focus solely on textual content. We introduce SPIQA (Scientific Paper Image Question Answering), the first large-scale QA dataset specifically designed to interpret complex figures and tables within the context of scientific research articles across various domains of computer science. Leveraging the breadth of expertise and ability of multimodal large language models (MLLMs) to understand figures, we employ automatic and manual curation to create the dataset. We craft an information-seeking task on interleaved images and text that involves multiple images covering plots, charts, tables, schematic diagrams, and result visualizations. SPIQA comprises 270K questions divided into training, validation, and three different evaluation splits. Through extensive experiments with 12 prominent foundational models, we evaluate the ability of current multimodal systems to comprehend the nuanced aspects of research articles. Additionally, we propose a Chain-of-Thought (CoT) evaluation strategy with in-context retrieval that allows fine-grained, step-by-step assessment and improves model performance. We further explore the upper bounds of performance enhancement with additional textual information, highlighting its promising potential for future research and the dataset’s impact on revolutionizing how we interact with scientific literature. equal technical contribution, work done as a student researcher at Google Research.lead, corresponding author (vsubhashini@google.com)
@@ 1 Introduction
Numerous datasets have been curated to evaluate the QA abilities of Large Language Models (LLMs) on various documents, including factual documents [85, 62, 22, 86, 32, 12, 68, 70], book chapters [28, 29, 51], news articles [74, 36] and more. However, understanding scientific papers poses unique challenges as the systems must comprehend underlying theoretical implications with domain-specific terminologies and verify claims with experimental results and visualizations. There are also several datasets that focus on the comprehension of standalone science diagrams [27, 24, 18, 65], mathematical figures [45, 87], charts [5, 48, 49], plots [50], tables [60, 76, 9, 8, 11, 26, 52] and retrieval [71]. However, simultaneously reasoning over all figures and associated text in a scientific article necessitates both multimodal and long-context capabilities.
In contemporary research, figures and tables with associated captions are crucial to understand the motivation and contributions of a work. Comprehending such multimodal content presents a significant real-world challenge. In this work, we introduce SPIQA (Scientific Paper Image Question Answering), the first large-scale QA dataset specifically designed to interpret complex figures and tables in the context of scientific research articles across various domains of computer science. We also develop three interleaved image-text tasks using the SPIQA dataset to assess growing long-context capabilities of MLLMs: (1) Direct QA with figures and tables, where the systems require to answer questions after seeing all figures and tables from a scientific paper, (2) Direct QA with full paper, where the systems analyze the whole paper text along with figures and tables to answer questions and (3) CoT QA, a chain of thought (CoT) retrieval based QA where systems need to first identify helpful figures and then answer the question, allowing to evaluate the models for fine-grained reasoning and grounding capability. We collect the PDFs and source TeXs of 26K papers from various domains of computer science published in top-tier academic conferences and generate 270K question-answer-rationale triplets focusing on the papers’ figures and tables.We further filter and augment two existing scientific QA datasets, QASA [35] and QASPER [14], to identify questions requiring reasoning over figures, tables, and textual paragraphs. Overall, SPIQA contains one training, one validation, and three evaluations split with varying difficulty levels, allowing us to fine-tune and assess the capabilities of large multimodal systems to understand complex scientific research papers. An illustration of the three proposed tasks with example prompts is presented in Figure 1.
@@ 2 Related Works
Full Text Figs & Tabs
BioAsq [75, 30] Human experts 3.2K -- PubMed Biomedical
BioRead [58] Cloze-style 16.4M 3.4M papers PubMed Biomedical
BioMRC [59] Cloze-style 812K 25M abstracts Pubtator Biomedical
emrQA [56] Cloze-style 455K 2.4K clinical notes i2b2 datasets EMRs
PubMedQA [21] Human experts 1K 120K abstracts PubMed Biomedical
BioASQ-QA [31] Human experts 4.7K -- PubMed Biomedical
ArgSciChat [66] Human Experts 41 dialogues 20 papers arXiv NLP
QASPER [14] Human experts 5K 1.5K papers S2ORC NLP
QASA [35] Human experts 1.8K 112 papers S2ORC AI/ML
Table 1: Comparison of SPIQA with existing scientific question answering datasets. SPIQA is a large-scale free-form QA corpus that focuses on the holistic aspects of scientific papers, including figures, tables, and full paper text. A dash (--) indicates the number is unreported.
@@
Question-answering on long documents is a challenging real-world task that has attracted increasing attention in recent years, following the success of the long-context reasoning ability of LLMs [41, 10, 72, 63, 54, 2, 37, 73, 16]. Though there exist numerous general-domain document QA datasets [85, 62, 22, 86, 32, 12, 68, 70, 79, 61, 84, 89, 78, 74, 36], understanding scientific papers requires domain-specific expertise and reasoning capability, and poses a more significant challenge.
@@
Datasets for QA on Scientific Papers. In the early days, cloze-style academic paper QA datasets [58, 56, 83, 69, 59] were automatically constructed by extracting entities and relations and matching them with structure knowledge resources. The questions in such datasets follow a pre-defined format; hence, they are unsuitable for real-world usage where the reader asks detailed open-ended questions [32]. To overcome these issues, PubMedQA [21], BioAsq [30] and QASPER [14] construct corpora of 1K, 3.2K, and 5K human-written questions, respectively. However, the annotators of these datasets only read the abstracts when writing the questions, and hence, most questions are simple and can be answerable in yes/no format or with short extractive spans. ArgSciChat [66] proposes a dataset of argumentive dialogues between scientists on 20 NLP papers. Closer to our work, QASA [35] generates 1798 free-form advanced questions on AI/ML papers where the annotators read the whole paper. However, QASA questions are answerable just from the text paragraphs, neglecting the structured information present in terms of figures and tables. Table 1 presents detailed comparisons.
@@
Datasets for QA on Scientific Diagrams. Solving mathematical problems in a visual context has emerged as a complex reasoning task for MLLMs. Prior attempts, such as GeoQA [7], UniGeo [6], and Geometry3K [46], have exclusively focused on solving geometry-oriented questions. In a different line of research, datasets like DVQA [23], LEAF-QA [5], LEAFQA++ [67], FigureQA [24], PlotQA [50], and ChartQA [48] have been constructed to solve plot and chart-oriented questions. Additionally, there are datasets for QA purely on tabular data, including WTQ [60], TableQA [76], SQA [20], HiTab [11], AIT-QA [26], FetaQA [52], MultiTabQA [55]. More recently, MathVista [45], MathVerse [87], and ArXivQA [38] have integrated different scientific diagrams to develop benchmarks with a wider variety of tasks. However, existing datasets focus on asking questions about standalone figures or tables (additional comparisons are in Appendix Sec. C). SPIQA bridges this gap by proposing a new scientific QA benchmark that includes questions that require simultaneous reasoning over figures, tables, and textual paragraphs, allowing a more integrated understanding of scientific documents.
Our proposed SPIQA dataset distinguishes itself from existing scientific question answering (QA) benchmarks in several significant ways. Firstly, it presents a large-scale, open-ended QA dataset to encompass diverse domains of computer science, providing a comprehensive evaluation framework for scientific QA systems. Secondly, the questions and answers in SPIQA necessitate an understanding of complex visual elements, including figures and tables, in conjunction with textual content and domain knowledge. This requirement simulates real-world scenarios where researchers must synthesize information from multiple sources to answer complex questions. Thirdly, we introduce a new Chain-of-Thought (CoT) QA paradigm in interleaved image-text comprehension, which involves a two-stage process wherein models first identify relevant figures and tables, and then generate answers based on this context. This step-by-step approach enables a more fine-grained assessment of the reasoning capabilities of baseline systems, providing valuable insights into their strengths and limitations.
@@ 3.1 Collection Guidelines and Task Formulation
  • CoT QA: To assess the step-by-step reasoning abilities of MLLMs, CoT QA requires systems first to identify relevant figures and tables and then answer the question. For simplicity, CoT QA does not involve full paper text. Detailed prompts for each task are included in the code.
3.2 SPIQA: Data Collection, Question Generation and Filtering
Automatic Question Generation and Filtering. After gathering the main text, figures, and tables from research papers, the next step is to generate high-quality question-answer pairs that cover all aspects of the articles. Manually annotating quality questions requires domain expertise and a deep understanding of the research papers, resulting in existing human-annotated datasets being small-scale or solely based on abstracts [75, 21, 30, 14, 35]. To bridge this gap, we automatically generate QAs by leveraging recent advances in powerful multi-modal large language models. We first conducted a pilot study by selecting 30 papers from various domains and experimenting with multiple models. In our approach, we presented one figure or table to the model, along with passages referencing the figure. We then asked the models to generate a question, answer, and rationale that requires a holistic understanding of the figure or table in the context of the paper. We manually verified the quality of the generated QAs by each model using the following criterion: $(i)$ Answering the question would require understanding of the figure or table. $(ii)$ The generated answer is correct and to the point. $(iii)$ The question is neither too trivial nor too specific to the figure or table. The questions looked promising and we proceeded to generate 270,194 QAs over 25,859 papers using Gemini 1.5 pro.
After generating QAs, we divide the dataset into three splits: 200 papers for evaluation, 200 for validation, and the remaining 25,459 for training, ensuring each split consists of papers from all domains. Despite the good question quality, we perform additional manual filtering on the evaluation set to exclude any inconsequential and incorrect QAs. In addition to the three criteria of the pilot study, the filtering guideline contains the following standards: $(iv)$ If two or more questions from a paper are similar, keep one. $(v)$ If the question is entirely based on the passage, discard the QA. $(vi)$ If the answer is not clear, e.g., the answer says ‘It is hard to answer the question based on the given information‘ or ‘The answer is not evident from the given passage‘, discard the QA. $(vii)$ If the QA includes phrases like ‘Based on the passage,’ modify it because we show all figures and tables to the model at once during evaluation. Appendix Sec. G includes the QA generation prompts, filtering guidelines and UI. We double annotated a small set of questions and found 88% agreement between annotators on which questions to discard. Overall, only around 11% of questions were discarded when filtering the evaluation set, and hence no additional filtering was performed on the train and validation splits. After filtering, the final evaluation set contains 666 questions, and we refer to this split as test-A.
@@ 3.2 SPIQA: Data Collection, Question Generation and Filtering
Additional Evaluation Sets with Human Written QAs. Since the questions and answers in test-A are automatically generated by Gemini, to develop a robust and complete QA benchmark on scientific articles, we curate two additional test sets with human-written QAs by utilizing two existing datasets, QASA [35] and QASPER [14]. However, the human annotators for these datasets mainly considered the paper text. We aim to identify the questions in these datasets that require reasoning with figures or tables for comprehensive answers. For every question, both QASA and QASPER provide evidence sentences from the paper, which are crucial for answering. We parsed these evidence sentences to look for phrases like ‘According to Figure x,’ ‘As shown in Table y,’ ‘Figure z presents,’ and separate such questions, and treat the corresponding figure and tables as evidence. We collect the PDFs for the papers in QASA and QASPER from arXiv, and extract the figures using PDFFigures 2.0. We conduct additional manual filtering to verify that the identified evidence figures help answer the questions. This process resulted in 228 questions from QASA and 493 questions from QASPER, where figures or tables are helpful for answering. We refer to these sets generated from QASA and QASPER as test-B and test-C, respectively. We additionally took a mix of human and model generated questions, and asked annotators to indicate whether each question was generated by a human, machine or if they’re unsure and 85% of the time, annotators indicated they were unsure, indicating the indistinguishability of LLM generated questions from human written ones.
@@ 3.3 Dataset Analysis
Splits & Statistics. The main statistics of the collected papers, generated questions, and data splits are presented in Tables 2 and 4. SPIQA encompasses 25,859 computer science papers published in top-tier conferences between 2018-2023, containing 152,487 figures and 117,707 tables. Figure 2 shows the distribution of figures and tables in every paper. We generate a total of 270,194 QAs focusing on the figures and tables along with the main text of the papers. The questions and answers, on average, contain 12.98 and 14.56 words, respectively. Notably, we observe high variances in their lengths - 20.47 for questions and 243.29 for answers - indicating a diverse range of patterns in SPIQA. The training, validation, and test-A splits include papers from every source conference and ensure that questions from the same paper remain in the same split. The test-B and test-C splits are generated from QASA [35] and QASPER [14] and contain human-written QAs. Examples from the dataset, illustrated in Figure 4, highlight two different questions centered on different types of figures. Appendix Sec. C provides a detailed comparison of SPIQA with existing scientific QA datasets.
@@
Split # Papers # Ques. # Figures # Tables
Sche. Plots & Charts Vis. Others.
Train 25,459 262,524 44,008 70,041 27,297 6,450 114,728
Val 200 2,085 360 582 173 55 915
test-A 118 666 154 301 131 95 434
test-B 65 228 147 156 133 17 341
test-C 314 493 415 404 26 66 1332
@@ 4 LLMLogScore (L3Score): An Improved Metric for Free-form QA
We alleviate the necessity of a predefined scale range and detailed interpretation of every point in the scale by proposing LLMLogScore (L3Score), which directly uses the log-likelihood probabilities generated by an LLM for evaluating the answers. We use GPT-4o [54] as our LLM, show the model the candidate and the ground-truth answers, and ask if the semantic meaning of the candidate and the ground-truth are similar in the context of the question. The LLM is expected to answer yes or no in a single word. The prompt used for calculating L3Score is in Appendix Sec. F. Next, instead of considering the final response by the LLM, we look into the top-5333We choose $n=5$ for top-$n$ log probabilities as this is the highest value of $n$ returned by the OpenAI API. log probabilities and corresponding tokens, and we define the L3Score metric as follows::::
@@
Figure 4 shows QA evaluations where L3Score is more appropriate than ROUGE and BERTScore. We provide additional examples comparing L3Score with other LLM-based metrics in Appendix H.